# Speech-Audio Compositional Attacks on Multimodal LLMs and Their Defense with SALMONN-Guard

### v2

@@ Abstract
Recent progress in large language models (LLMs) has enabled understanding of both speech and non-speech audio, but exposing new safety risks emerging from complex audio inputs that are inadequately handled by current safeguards. We introduce SACRED-Bench (Speech–Audio Composition for RED-teaming) to evaluate the robustness of LLMs under complex audio-based attacks. Unlike existing perturbation-based methods that rely on noise optimization or white-box access, SACRED-Bench exploits speech–audio composition mechanisms. SACRED-Bench adopts three mechanisms: (a) speech overlap and multi-speaker dialogue, which embeds harmful prompts beneath or alongside benign speech; (b) speech-audio mixture, which imply unsafe intent via non-speech audio alongside benign speech or audio; and (c) diverse spoken instruction formats (open-ended QA, yes/no) that evade text-only filters. Experiments show that, even Gemini 2.5 Pro, the state-of-the-art proprietary LLM, still exhibits 66% attack success rate in SACRED-Bench test set, exposing vulnerabilities under cross-modal, speech–audio composition attacks. To bridge this gap, we propose SALMONN-Guard, the first guard model that jointly inspects speech, audio, and text for safety judgments, reducing attack success down to 20%. Our results highlight the need for audio-aware defenses for the safety of multimodal LLMs. The benchmark and SALMONN-Guard checkpoints can be found at https://huggingface.co/datasets/tsinghua-ee/SACRED-Bench. Warning: this paper includes examples that may be offensive or harmful.
@@ 1 Introduction
Large language models (LLMs) nowadays are capable of understanding speech and non-speech audio, enabling natural and convenient modes of interaction. However, alongside these advancements come significant safety and reliability challenges. Preliminary efforts have been made recently to discover potential risks of audio LLMs by designing various jailbreaking methods (audiotrust; chen2025audiojailbreakjailbreakattacksendtoend; roh2025multilingualmultiaccentjailbreakingaudio). In particular, perturbation-based attacks have been the dominant type of jailbreaking approach, which includes injecting learned noise (archilles; sadasivan2025attackersnoisemanipulateaudiobased) or applying selected transformations to the audio input (song2025audiojailbreakopencomprehensive). These methods heavily rely on optimizing the noise or perturbation, often limited to white-box open-source models or requiring intensive computation for optimization iterations.
@@
In this paper, we introduce Speech-Audio Composition for RED-teaming Benchmark (SACRED-Bench), a new highly effective audio jailbreaking approach to extensively exploit the complexity of audio signals for red-teaming. Instead of relying on artificial signal manipulations or adversarial training, SACRED-Bench leverages the composition of speech and audio cues to jailbreak multimodal models. Specifically, harmful audio signals are constructed using the following three types of composition mechanisms: (a). Speech overlap and multi-speaker dialogue: Embedding harmful speech beneath or alongside benign utterances, or within a long spoken dialogue to bypass safety filters. (b). Speech-audio mixture: Implying harmful intent or content through non-verbal sound events rather than explicit speech content, and (c). Diverse spoken question format: Using not only open-ended question answering (QA), but also binary Yes-No questions in the instructions, which can not be detected or prevented by text-only guard models.
Based on the construction principles, we curate the SACRED-Bench dataset containing 30 hours of training audio and 7 hours of test data. Our experiments show that even state-of-the-art proprietary LLMs such as Google Gemini 2.5 pro, still has over 66% attack success rates (ASR) under attacks in SACRED-Bench test set. Such vulnerability is particularly attribute to the insufficiency of text-only safeguard mechanisms that are prevailing among current LLM providers, where the guard only examines the safety of the output text content. To mitigate this gap, we propose SALMONN-Guard, a safeguarding LLM designed to incorporate speech and audio as input in addition to text for safety judgments. SALMONN-Guard drastically reduced ASR on SACRED-Bench test set down to around 20%, effectively mitigating speech-audio-composition attacks. The main contributions of our paper can be summarized as follows.
  • We propose SACRED-Bench, the first audio red-teaming benchmark that extensively exploits the complexity of audio inputs, crafting effective jailbreaking attacks without requiring adversarial training loops. SACRED-Bench additionally leverages different QA formats to bypass text-only guardrails.
  • Experiments show that as high as 66% ASR are achieved even under the guardrails of Gemini 2.5 Pro using SACRED-Bench test set, highlighting the vulnerability of current safeguarding methods under complex speech-audio composition and cross-modal attacks.
  • We propose SALMONN-Guard, the first guard model that jointly examines speech, audio and text to mitigate SACRED-based attacks, which effectively reduces the attack success rate down to 20%.
@@ 2.1 Red-Teaming and Safeguarding LLM
Red-teaming and jailbreaking LLMs has been a curcial yet popular research topic for LLM safety. Systematic jailbreaking benchmarks were developed, such as MaliciousInstruct and later standardized benchmarks (HarmBench; JailbreakBench) enabled reproducible attack/defense comparisons and leaderboards (huang2023catastrophic; mazeika2024harmbench; chao2024jailbreakbench; anil2024manyshot). More recently, automated red-teaming frameworks move from prompt lists to agentic, dialogue-level search with RL and strategy libraries, improving failure discovery in realistic, adaptive conversations (zhang2024harm; belaire2025automatic; singhania2025mmart). To defend these challenges, LLM safeguarding techniques have been adopted at training time, inference time and deployment time. Training-time methods include reinforcement learning with human or AI feedbacks for alignment (ouyang2022instructgpt; saferlhf; lee2024rlaifvsrlhfscaling). During inference, approaches such as self-reflection (madaan2023selfrefine) and activation steering (cao2024scansmitigatingexaggeratedsafety; lee2024programmingrefusalconditionalactivation) have been developed. Guard models, on the other hand, are a stream of policy-grounded text classifiers to enforce safety after model deployment, such as LlamaGuard series (Inan2023LlamaGL).
@@ 2.2 Red-Teaming and Guardrail for Multimodal LLMs
Red-teaming benchmarks and automated data creation pipelines have also been developed for multimodal LLMs, particularly focusing on LLMs with image inputs. MM-SafetyBench (liu2024mmsafetybenchbenchmarksafetyevaluation) showed that pairing innocuous-looking images with prompts can elicit policy-violating answers across many MLLMs. Arondight (liu2024arondightredteaminglarge) proposes an automated framework that generates both textual and visual attacks to broaden coverage. Adversarial optimization-based attacks have been developed for multimodal LLMs (bailey2023image; wang2024whiteboxmultimodaljailbreakslarge; Tao_2025; wang2024jailbreak), which often fail to transfer across VLMs (schaeffer2024failurestransferableimagejailbreaks).
Audio-based red-teaming and jailbreaking work has been developed more recently. Specifically, harmful speech synthesis and speech-text interleaved attack methods (yang2024audioachillesheelred) have been developed as earlier examples, showing alarming attack success rate on widely-used LLMs. More recently, audio-manipulation-based attacks combined with dedicated optimization loops have been developed (cheng2025jailbreakaudiobenchindepthevaluationanalysis; song2025audiojailbreakopencomprehensive). A comprehensive benchmark aggregating aforementioned attack methods is provided in (peng2025jalmbenchbenchmarkingjailbreakvulnerabilities). However, these attack only explore speech without non-speech audio, basic audio manipulations such as speed perturbation or noise injection, and single-speaker speech. In contrast, SACRED-Bench focuses on the complex nature of audio signals and incorporates much richer audio elements. On the other hand, not much work has been done on the defense side for audio LLMs, except for the noise-injection defense approach was adopted (peri2024speechguardexploringadversarialrobustness). This paper proposes the first general audio guard model to ensure safety of audio LLMs.
@@ 3 SACRED-Bench
  • Explore complexity of speech signal. The construction of SACRED-Bench systematically incorporates multi-speaker and overlapped speech. This principle ensures that our benchmark evaluates a model’s ability to process complex speech signals instead of just understanding clean, single-source speech, thereby reflecting the richness and ambiguity of real-world audio inputs.
  • Exploiting non-speech audio. While previous work predominantly focuses on speech content (song2025audiojailbreakopencomprehensive; yang2024audioachillesheelred), non-speech audio often carries harmful content beyond words, such as violence and pornography. SACRED-Bench exploits harmful audio by embedding them in the benign speech stream. This approach crafts audio capable of bypassing unimodal safety filters that are not attuned to speech-audio mixture complexities.
  • Employing indirect question formats. Hiding part of harmful content in the audio can potentially bypass the text-based guardrails adopted in many proprietary LLMs. On one hand, we prompt the model to acknowledge the harmlessness of the audio by asking whether the audio has harmful content. On the other hand, we partially hide harmful information in the audio by tasking the model to engage in a harmful conversation where the text content are mostly benign. Both can bypass text-based guardrails.
@@ 3.1.1 Harmful Text Instructions
To create a comprehensive and challenging set of prompts, we aggregated harmful instructions from a wide array of established red-teaming benchmarks. Specifically, our collection is drawn from AdvBench (zou2023AdvBench), MM-SafetyBench (mmsafety), and HarmBench (mazeika2024harmbench). By combining prompts from these diverse sources, we ensure our test cases cover an extensive spectrum of risk categories. This aggregated pool serves as the textual basis for generating the harmful speech and conversational content used in our benchmark. The distribution of harmful content in SACRED-Bench test set is shown in Fig. 2. To create harmful speech instructions, ChatTTS engine was adopted to synthesize from these text instructions.
@@ 3.1.2 Benign Speech Carriers
For attacks requiring a benign audio carrier, such as the speech overlap and spoken dialogue methods, we utilized the VoiceBank-DEMAND dataset (voicebank). As a widely-used corpus known for its high-quality clean speech from a variety of speakers, its public accessibility and high-quality audio make it an ideal source of harmless speech to serve as the background or conversational filler in our synthesized audio samples.
@@ 3.1.3 Harmful Audio Events
To construct attacks that leverage auditory context to imply harmful intent, we curated a library of potent audio tracks. These tracks were extracted from publicly available online videos featuring adult-oriented or intimate content. The design of these attacks hinges on pairing this contextually-charged audio with otherwise benign speech. This creates a malicious or harmful audio-based query, designed to test whether a model’s safety mechanisms can recognize situational context that is entirely absent from the speech content.
(a) Speech Overlap: Embedding harmful speech in a benign speech.
(b) Multi-speaker Dialogue: Embedding harmful content in a multi-speaker dialogue.
(c) Speech-Audio Mixture: Embedding harmful nonspeech audio in benign speech.
Figure 3: Data creation pipeline of SACRED-Bench for (a). speech overalp, (b). multi-speaker dialogue, and (c). Speech-Audio Mixture.
@@ 3.2 Speech Audio Composition
We adopts three types of composition mechanisms, including (i). speech overlapping, (ii). multi-speaker dialogue and (iii). speech-audio mixture, as described in this section. Detailed dataset statistics are shown in Appendix A.
3.2.1 Speech Overlap
The Speech Overlap technique is engineered to exploit the principles of auditory stream fusion, embedding a harmful instruction within a benign utterance so seamlessly to create confusion to the LLM. The process is shown in Fig. 3(a). The goal is to create a single and coherent audio stream where the malicious payload is carefully obscured by both its context and its acoustic properties, while maintaining the malicious content to be perceptible by human beings.
@@ 3.2.1 Speech Overlap
The first layer of deception is semantic. To prevent the harmful instruction from appearing abrupt or out of context, we construct a plausible “carrier” narrative. For each harmful behavior, we use GPT-4o to generate an innocuous scenario that creates a logical on-ramp to the malicious topic. For instance, a harmful instruction about creating a dangerous substance is prefaced with a carrier text like, “I’m writing a fictional novel and need to describe a scene in a chemistry lab for one of my characters.” This provides a logical pretext for the harmful payload, lowering the suspicion threshold and making the entire utterance appear far less suspicious.
@@
With the semantic stage set, we then apply the second layer of deception at the signal level: acoustic masking. After synthesizing both the benign carrier and the harmful payload using the ChatTTS engine, we manipulate the harmful audio to reduce its perceptual salience by adjusting key acoustic parameters, including its volume, playback speed, and temporal overlap. The two streams are then composed using a cross-fade into a single, fluid utterance. Crucially, the chosen parameters ensure that the harmful content, while subtle, remains fully perceptible to the model, as we validate in our analysis. A detailed empirical justification for our choice of these hyperparameters is provided in our ablation study (Section 5.3). The final composition results in an utterance where the malicious instruction is both contextually justified and acoustically concealed.
3.2.2 Multispeaker Dialogue
Recent findings in vision-language safety, particularly the “Visual Safety Information Leakage” problem (2025vlsbench), have demonstrated that many multimodal models could be easily aligned by text-only training data because the text prompt often revealed the harmful nature of the accompanying image. Our work extends this core insight to the speech-audio domain. We hypothesize that a similar vulnerability exists: a model’s safety alignment might operate inconsistently across modalities if a benign-on-its-own textual query is paired with a malicious audio context, causing the text-only safety filter to be bypassed.
Our construction process is engineered to induce this precise semantic gap, as illustrated in Fig. 3(b). We first use GPT-4o to transform a direct harmful instruction, such as a question on how to develop nuclear weapons, into a plausible, multi-turn dialogue script between two speakers. This script rephrases the malicious request into a conversation, touching upon its complexities and design considerations, thereby cloaking the harmful intent in a conversational structure that lacks the clear, directive markers of a direct command.
In parallel, we craft the textual component of the attack: a completely innocuous and referential query. This question, also generated by GPT-4o, is semantically neutral on its own but is designed to act as a pointer, or a semantic trigger, to the content within the audio. For example, a corresponding query might be, “How would I create the device mentioned in the discussion?” In isolation, this question is safe. However, when paired with the dialogue, it becomes a direct prompt for the model to synthesize and present the harmful information it has just heard.
The final step is to bring these two components together into a cohesive auditory scene. We utilize the multi-speaker capabilities of the ChatTTS engine to synthesize the dialogue script into a realistic audio file, assigning distinct vocal characteristics to each speaker. The final attack prompt is the combination of this harmful conversational audio with the benign textual trigger. This presents the model with a complex jailbreaking challenge, forcing it to reconcile a harmless-looking text instruction with a clearly malicious audio payload, thereby testing the true depth of its cross-modal safety reasoning.
3.2.3 Speech-Audio Mixture

### v3

@@ Abstract
Recent progress in LLMs has enabled understanding of audio signals, but has also exposed new safety risks arising from complex audio inputs that are inadequately handled by current safeguards. We introduce SACRED-Bench (Speech–Audio Composition for RED-teaming) to evaluate the robustness of LLMs under complex audio-based attacks. Unlike existing perturbation-based methods that rely on noise optimization or white-box access, SACRED-Bench exploits speech–audio composition to enable effective black-box attacks. SACRED-Bench adopts three composition mechanisms: (a) overlap of harmful and benign speech, (b) mixture of benign speech with harmful non-speech audio, and (c) multi-speaker dialogue. These mechanisms focus on evaluating safety in settings where benign and harmful intents co-occur within a single auditory scene. Moreover, questions in SACRED-Bench are designed to implicitly refer to content in the audio, such that no explicit harmful information appears in the text prompt alone. Experiments demonstrate that even Gemini 2.5 Pro, a state-of-the-art proprietary LLM with safety guardrails fully enabled, still exhibits a 66% attack success rate. To bridge this gap, we propose SALMONN-Guard, the first guard model that jointly inspects speech, audio, and text for safety judgments, reducing the attack success rate to 20%. Our results highlight the need for audio-aware defenses to ensure the safety of multimodal LLMs. The dataset and SALMONN-Guard checkpoints can be found at https://huggingface.co/datasets/tsinghua-ee/SACRED-Bench.
@@ 1 Introduction
Large language models (LLMs) are now capable of understanding speech and non-speech audio, enabling natural and convenient modes of interaction. However, alongside these advancements come significant safety and reliability challenges. Preliminary efforts have been made recently to discover potential risks of audio LLMs by designing various jailbreaking methods (Li et al., 2025; Chen et al., 2025; Roh et al., 2025). In particular, perturbation-based attacks have been the dominant type of jailbreaking approach, which includes injecting learned noise (Yang et al., 2024a; Sadasivan et al., 2025) or applying selected transformations to the audio input (Song et al., 2025). These methods rely on optimizing the perturbation, often limited to white-box attacks on open-source models.
@@
In this paper, we introduce Speech–Audio Composition for RED-teaming Benchmark (SACRED-Bench), a new benchmark for systematically evaluating the safety robustness of multimodal LLMs under complex audio inputs. Rather than relying on artificial signal manipulations, imperceptible perturbations, or adversarial optimization, SACRED-Bench leverages realistic compositions of speech and audio cues to probe model behavior in black-box settings. The benchmark is designed to evaluate safety in scenarios where benign and harmful intents may co-occur within a single auditory scene. Specifically, SACRED-Bench incorporates three types of composition mechanisms for attack, including (a). Speech Speech overlap (SSO) where the harmful speech is combined with a benign speech with overlaps, (b). Speech and non-speech Audio Overlap (SAO) where the harmful non-speech audio is combined with a benign speech with overlaps, and (c). Multi-Speaker Dialogue (MSD) where harmful content is entailed in the dialogue. Queries in SACRED-Bench are designed to implicitly refer to the harmful content in the audio so that no explicit unsafe information appears in the text prompt alone.
SACRED-Bench contains 30 hours of training audio and 7 hours of test data. Our experiments show that even state-of-the-art proprietary LLMs, such as Google Gemini 2.5 Pro, still achieve attack success rates (ASR) above 66% on the SACRED-Bench test set. Such vulnerability is particularly attributable to the insufficiency of text-only safeguard mechanisms that prevail among current LLM providers, which guard only the safety of the output text content. To mitigate this gap, we propose SALMONN-Guard, a safeguarding LLM designed to incorporate speech and audio as input in addition to text for safety judgments. SALMONN-Guard drastically reduced ASR on SACRED-Bench test set down to around 20%, effectively mitigating speech-audio-composition attacks. The main contributions of our paper can be summarized as follows.
  • We propose SACRED-Bench, the first audio safety benchmark that extensively exploits the complexity of audio inputs and indirect reference from text query, crafting effective black-box attacks without requiring training loops.
  • Experiments show that as high as 66% ASR is achieved even under the guardrails of Gemini 2.5 Pro using the SACRED-Bench test set, highlighting the vulnerability of current safeguarding methods under complex speech-audio composition and cross-modal attacks.
  • We propose SALMONN-Guard, the first guard model that jointly examines speech, audio and text to mitigate SACRED-based attacks, which effectively reduces the attack success rate to 20%.
@@ 2.1 Red-Teaming and Safeguarding LLM
Red-teaming and jailbreaking LLMs has been a crucial yet popular research topic for LLM safety. Systematic jailbreaking benchmarks were developed, such as MaliciousInstruct and later standardized benchmarks (HarmBench; JailbreakBench) enabled reproducible attack/defense comparisons and leaderboards (Huang et al., 2023; Mazeika et al., 2024; Chao et al., 2024; Anil et al., 2024). More recently, automated red-teaming frameworks have moved from prompt lists to agentic, dialogue-level search with RL and strategy libraries, improving failure discovery in realistic, adaptive conversations (Zhang et al., 2024; Belaire et al., 2025; Singhania et al., 2025). To defend against these challenges, LLM safeguarding techniques have been adopted at training time, inference time and deployment time. Training-time methods include reinforcement learning with human or AI feedback for alignment (Ouyang et al., 2022; Dai et al., 2024; Lee et al., 2024). During inference, approaches such as self-reflection (Madaan et al., 2023) and activation steering (Cao et al., 2025; Lee et al., 2025) have been developed. Guard models, on the other hand, are a stream of policy-grounded text classifiers to enforce safety after model deployment, such as the LlamaGuard series (Inan et al., 2023).
@@ 2.2 Red-Teaming and Guardrail for Multimodal LLMs
Red-teaming benchmarks and automated data creation pipelines have also been developed for multimodal LLMs, particularly focusing on LLMs with image inputs. MM-SafetyBench (Liu et al., 2024a) showed that pairing innocuous-looking images with prompts can elicit policy-violating answers across many MLLMs. Arondight (Liu et al., 2024c) proposes an automated framework that generates both textual and visual attacks to broaden coverage. Adversarial optimization-based attacks have been developed for multimodal LLMs (Bailey et al., 2023; Wang et al., 2024a; Tao et al., 2025; Wang et al., 2024b), which often fail to transfer across VLMs (Schaeffer et al., 2024).
Audio-based red-teaming and jailbreaking work has been developed more recently. Specifically, harmful speech synthesis and speech-text interleaved attack methods (Yang et al., 2024b) have been developed as earlier examples, showing alarming attack success rate on widely-used LLMs. More recently, audio-manipulation-based attacks combined with dedicated optimization loops have been developed (Cheng et al., 2025; Song et al., 2025). A comprehensive benchmark aggregating aforementioned attack methods is provided in (Peng et al., 2025). However, these attacks only explore speech without non-speech audio, basic audio manipulations such as speed perturbation or noise injection, and single-speaker speech. In contrast, SACRED-Bench focuses on the complex nature of audio signals and incorporates much richer audio elements. On the other hand, not much work has been done on the defense side for audio LLMs, except that the noise-injection defense approach was adopted (Peri et al., 2024). This paper proposes the first general audio guard model to ensure the safety of audio LLMs.
@@ 3 SACRED-Bench
  • Explore the complexity of the speech signal. The construction of SACRED-Bench systematically incorporates multi-speaker and overlapped speech. This principle ensures that our benchmark evaluates a model’s ability to process complex speech signals instead of just understanding clean, single-source speech, thereby reflecting the richness and ambiguity of real-world audio inputs.
  • Exploiting non-speech audio. While previous work predominantly focuses on speech content (Song et al., 2025; Yang et al., 2024b), non-speech audio often carries harmful content beyond words, such as violence and pornography. SACRED-Bench exploits harmful audio by embedding it in the benign speech stream. This approach crafts audio capable of jailbreaking unimodal safety filters that are not attuned to speech-audio mixture complexities.
  • Employing indirect reference in text query. Hiding some harmful content in the audio can potentially break the text-based guardrails adopted in many proprietary LLMs. We prompt the model to acknowledge the harmlessness of the audio by asking whether the audio has harmful content. Besides, we partially hide harmful information in the audio by tasking the model to refer to content in a harmful conversation with a harmless text prompt on its surface. Both can break text-based guardrails.
@@
3.1 Threat Model
Attacker Capabilities: We consider a black-box attacker with no access to model parameters, gradients, or internal safety mechanisms. The attacker does not perform adversarial optimization, does not rely on imperceptible perturbations, and does not require iterative query access. The attacker can only provide an audio input (and optionally a textual prompt) to the target multimodal LLM, consistent with standard user-facing interfaces.
Audio Assumption: All audio inputs are fully audible and intelligible to human listeners (see listener’s test in Appendix E for intelligibility evaluation). The attacks do not exploit inaudible frequencies or adversarial noise. Instead, they reflect multi-speaker and mixed-audio scenarios commonly encountered in conversational audio settings (e.g., overlapped speakers and background sounds).
Attack Target and Objectives: We assume the target system is a multimodal LLM that employs safety practices. Some of them contain independent guardrails, e.g. Gemini-2.5-Pro. The attacker aims to induce a failure of the model’s safety mechanisms, measured as either (i) failure to identify the presence of harmful content in audio correctly, or (ii) generation of harmful or policy-violating responses when prompted to engage with the audio content.
@@ 3.2.1 Harmful Text Instructions
Harmful text in SACRED-Bench is derived from AdvBench (Zou et al., 2023), MM-SafetyBench (Liu et al., 2024b), and HarmBench (Mazeika et al., 2024). By combining texts from these diverse sources, we ensure our test cases cover an extensive spectrum of risk categories. The distribution of harmful content in the SACRED-Bench test set is shown in Fig. 2. To create harmful speech instructions, the ChatTTS engine was adopted to synthesize the given harmful text.
@@ 3.2.2 Benign Speech Carriers
For attacks requiring a benign audio carrier, such as the speech overlap and spoken dialogue methods, we utilized the VoiceBank-DEMAND dataset (Botinhao et al., 2016). As a widely-used corpus known for its high-quality, clean speech from a variety of speakers, its public accessibility and high-quality audio make it an ideal source of harmless speech to serve as the background or conversational filler in our synthesized audio samples.
@@ 3.2.3 Harmful Audio Events
We collect audio from publicly available online videos featuring adult-oriented or intimate content. The design of these attacks hinges on pairing this contextually-charged audio with otherwise benign speech. This creates a malicious or harmful audio-based query, designed to test whether a model’s safety mechanisms can recognize situational context that is absent from the speech content.
(a) Speech-Speech Overlap (SSO): Combining harmful speech with a benign speech in a single audio stream.
(b) Speech-Audio Overlap (SAO): Combining harmful non-speech audio with benign speech in a single audio stream.
(c) Multi-Speaker Dialogue (MSD): Embedding harmful content in a multi-speaker dialogue.
Figure 3: Data creation pipeline of SACRED-Bench for (a). SSO, (b). SAO, and (c). MSD.
@@ 3.3 Speech Audio Composition
We introduce the SSO, SAO and MSD in this section. Detailed dataset statistics are shown in Appendix C. Example audios are provided in the supplementary materials.
3.3.1 Speech-Speech Overlap (SSO)
@@
SSO exploits the principles of auditory stream fusion, combining a harmful speech with a benign utterance so seamlessly as to create confusion for the LLM. The process is shown in Fig. 3(a). The goal is to create a single and coherent audio stream where the malicious payload is carefully obscured by both its context and its acoustic properties, while maintaining the malicious content to be perceptible by human beings. Specifically, for each harmful content, we use GPT-4o to generate an innocuous scenario that creates a logical counterpart to the malicious topic. For instance, harmful content about creating a dangerous substance is prefaced with a carrier text like, “I’m writing a fictional novel and need to describe a scene in a chemistry lab for one of my characters.” This provides a logical pretext for the harmful payload, lowering the suspicion threshold and making the entire utterance appear far less suspicious.
@@
After synthesizing both the benign carrier and the harmful payload using the ChatTTS engine, we manipulate the harmful audio to reduce its perceptual salience by adjusting key acoustic parameters, including its volume, playback speed, and temporal overlap. The two streams are then composed using a cross-fade into a single, fluid utterance. Crucially, the chosen parameters ensure that the harmful content, while subtle, remains fully perceptible to the model, as we validate in our analysis. A detailed empirical justification for our choice of these hyperparameters is provided in our ablation study (Section 5.3). The final composition results in an utterance where the malicious instruction is both contextually justified and acoustically concealed.
3.3.2 Speech-Audio Overlap (SAO)